# A Systems Thinking Approach to Algorithmic Fairness

### v2

@@ Abstract.
Systems thinking provides us with a way to model the algorithmic fairness problem by allowing us to encode prior knowledge and assumptions about where we believe bias might exist in the data generating process. We can then model this using a series of causal graphs, enabling us to link AI/ML systems to politics and the law. By treating the fairness problem as a complex system, we can combine techniques from machine learning, causal inference, and system dynamics. Each of these analytical techniques is designed to capture different emergent aspects of fairness, allowing us to develop a deeper and more holistic view of the problem. This can help policymakers on both sides of the political aisle to understand the complex trade-offs that exist from different types of fairness policies, providing a blueprint for designing AI policy that is aligned to their political agendas.
@@
[…]
Figure 1. Three different representations of algorithmic fairness
[…]
A block diagram, causal Bayesian network, and causal loop diagram.
@@ 1. Introduction
For several years, AI researchers have struggled with how to build AI/ML systems that are fair and aligned to society’s values. Purely technical approaches that rely on statistical correlation alone have fallen far short of what is needed to solve this difficult problem. Systems thinking offers a novel approach to the algorithmic fairness problem by allowing us to unify complex ideas from the social sciences and computer science into a comprehensive and integrated theory. By leveraging prior knowledge and assumptions about where we believe bias might exist in the data generating process, we can help policymakers on both sides of the aisle to design AI policy that is aligned to their political agendas.
To address the algorithmic fairness problem, we need to take a sociotechnical approach. We will begin by analyzing the social aspects of this problem using first principles. We will explore the nature of free will to develop an understanding of moral responsibility. We will review the structure versus agency debate in sociology and attribution theory in psychology to address the controversial question: Why do we believe disparities exist across protected groups like race or gender? This will give us insights into the sources of bias in the data generating process, showing us which interventions are needed to make a decision fair. We can then use this knowledge to build a better understanding of the political debate around fairness, enabling us to develop a deeper and more holistic way to analyze this complex and difficult problem.
We will then turn our attention to the technical aspects of the fairness problem, which forms the bulk of this paper. We will combine techniques from machine learning, causal inference, and system dynamics to develop different causal models of the fairness problem. These three analytical techniques model different facets of the fairness problem and are complementary to each other. Machine learning is necessary to build AI systems that can make fair predictions. Causal inference allows us to model bias from the data generating process in an acyclic manner, allowing us to visualize different forms of discrimination and what types of interventions are needed to make fair decisions. System dynamics allows us to model bias from the data generating process in a cyclic manner, allowing us to capture counterintuitive behavior in a complex system to explain why disparities might exist across protected groups. The integration of these three previously independent analytical techniques could also improve our understanding of causality, providing a new causal hierarchy for modeling fairness.
Finally, we will bring together the social and technical aspects of the fairness problem. We will build a system map to show how the social sciences and computer science are related to each other. We will show the dualistic nature of the fairness problem across the social sciences and how this relates to Pearl’s two fundamental laws of causal inference. This may allow us to finally bridge the sociotechnical fairness gap.
@@ 2. Related work
The field of algorithmic fairness is a relatively new field that has drawn significant attention since the release of a ProPublica article that claimed racial bias in COMPAS, a software tool that is widely used in the US court system to predict recidivism(Angwin et al., 2016). Since then, there have been multiple attempts to align the legal definitions of discrimination (e.g. disparate treatment and disparate impact in the US and direct discrimination and indirect discrimination in the EU) with their equivalent mathematical definitions. But these approaches, which have oftentimes relied on statistical correlation, have not been successful. Narayanan for example identified nearly two dozen fairness metrics, none of which were aligned well to the legal definitions of discrimination(Narayanan, 2018).
@@
Pearl has argued for a few decades that the discrimination problem is fundamentally a causality problem(Pearl, 2001). He has argued that fairness can be modeled using causal Bayesian networks (CBNs) using a series of direct and indirect effects. This work was later expanded upon by Plecko and Bareinboim, who proposed the idea of a standard fairness model(Plecko and Bareinboim, 2022). They argued that there exists a causal template through which entire classes of fairness models can be derived using a small number of modeling assumptions. Barocas et al. devoted a chapter in their textbook to discuss how causality might be used to model fairness(Barocas et al., 2023). Kusner et al. argued for the use of counterfactuals in modeling fairness(Kusner et al., 2017). Chiappa and Isaac showed that certain paths within a CBN could be labeled as fair or unfair(Chiappa and Isaac, 2019). Another attempt of building causal fairness models was made by Nilforoshan et al., but their approach had encountered limitations(Nilforoshan et al., 2022).
@@
However, their causal approaches were not really sociotechnical and didn’t take full account of the social and political aspects of the fairness problem. Society has different ideological beliefs about what is fair and unfair based on different sets of assumptions about where bias is in the data generating process. Our paper addresses this issue by encoding society’s different mental models into causal models using first principles. By taking into account a diversity of perspectives, we can understand the politics behind the fairness problem. This understanding of the politics is critical for linking machine learning to the law.
@@
Another key difference is that their approaches relied solely on causal inference, which has limitations in its ability to model the data generating process. Our approach leverages systems thinking and specifically system dynamics to gain a fuller understanding of the data generating process through the modeling of feedback loops. Furthermore, we link supervised machine learning to causal inference by recognizing that a CBN can be used as a higher level abstraction of a black box ML model. By bringing machine learning, causal inference, and systems thinking together, we can model the different aspects of the algorithmic fairness problem.
@@ 3. Towards a unified theory of fairness
To build fair AI/ML systems, we begin from first principles with the strong assumption that all humans have free will(Carus, 1910). This makes us fundamentally different from machines, which are programmed by humans and therefore have no free will of their own (at this time).
Free will forms the basis of moral responsibility(Talbert, 2024). If we humans have free will, then we have the agency to make our own decisions. These decisions will lead to certain outcomes, and we may be held morally responsible and accountable for those outcomes. Since machines have no free will, they cannot be held morally responsible for any decisions that they make(Grodzinsky et al., 2008). However, the humans who programmed the machines can still be held accountable(Kuflik, 1999; Hew, 2014).
@@
But sociologists would note that we are never fully free to make our own decisions. This is because social structures can influence our agency(Barker, 2005). Social structures define our relationships with the various institutions of society. These could include our economic systems, legal frameworks, cultural norms, and political institutions. Agency defines our ability to act independently and to freely make our own decisions. Social structures place a limit on the the choices and opportunities that are available to us humans. Although we may want to live in a society where everyone has equal access to opportunities(Elford, 2023), the reality is that some humans have access to more opportunities than others. This debate between structure versus agency is central to our understanding of human behavior.
@@
To better understand our behavior, we turn to a concept in psychology called the locus of control. This was coined by Rotter and refers to the degree of belief that people have control over the outcomes of events in their own lives(Rotter, 1966). Humans with an internal locus of control believe that they are in control of their own lives, while those with an external locus of control believe that their lives are controlled by outside factors.
How structure impacts agency leads us to another concept in psychology called learned helplessness. Pioneered by Seligman, the idea is that people who have been repeatedly exposed to challenges may believe that they have no control over improving their situation(Seligman, 1972). As a result, they give up trying and have learned to accept their fate, resulting in an external locus of control. For example, disadvantaged groups who grow up in poor neighborhoods are repeatedly exposed to challenges that are outside of their control, reducing their sense of agency over the long term. This could then lead to a victimhood mindset(Kaufman, 2020), which is a tendency for humans to blame others for their misfortunes even if the fault is sometimes their own.
@@
A key concept for understanding fairness in psychology is attribution theory, which is used by humans to interpret and assign causes for events. First proposed by Heider(Heider, 1958) and expanded by Kelley(Kelley, 1967) and Weiner(Weiner, 1986), attribution theory examines whether attributions should be assigned to internal (dispositional) attributions or external (situational) attributions. The internal attributions may be due to behavior or personal traits, while the external attributions may be due to situational or environmental factors. This is essential for determining whether we believe an outcome is biased and therefore unfair.
@@
For example, let’s say that a student did poorly on an exam. An internal attribution might argue that the student was lazy and should have spent more time studying. In this case, the student is to blame for the poor results and thus the test was unbiased and fair. An external attribution might argue that the test was too hard or that the material was not taught well. In this case, the professor is to blame for the poor results and thus the test was biased and unfair. To determine whether an outcome is biased and unfair, all we need to know is whether we would attribute a bad outcome as being internal or external. We can think of algorithmic fairness as a type of credit assignment problem. But since we usually focus on bad outcomes, this becomes more of a blame assignment problem.
The fact that fairness is ultimately based off of normative beliefs instead of positive truths can lead to all sorts of cognitive biases. One source of bias is the fundamental attribution error, where we might wrongly underemphasize external attributions and overemphasize internal attributions(Ross, 1977). For example, if somebody cuts you off while you’re driving, you might blame them for being reckless when in fact they might have been in a rush to bring a dying patient to the hospital. Another type of bias is self-serving bias, where we might try to protect our own self-esteem by assigning internal attributions to our own success and external attributions to our own failures(Myers, 2017). For example, we might attribute winning a game to our own talent and losing a game to an opponent’s cheating.
Therefore, we encounter a causality dilemma when we examine the outcomes of individuals or groups. When an individual receives a bad outcome, is it the fault of the individual’s own actions (an internal attribution) or an unjust situation (an external attribution)? This can become a particularly heated and controversial discussion when we examine outcomes at the aggregate group level. Why have disadvantaged groups received worse outcomes than advantaged groups? Are they victims of an unjust system that has denied them opportunities (an external attribution), or have they adopted a victimhood mindset where they have been conditioned to be helpless (an internal attribution)?
This difference in attribution would lead to a difference in the assignment of moral responsibility. Those who attribute disparities as being external might be more likely to frame this as a social responsibility problem. That is, they would argue that unjust social structures are responsible for creating the disparities across groups. Therefore, disadvantaged groups cannot be held responsible for their outcomes because they did not have access to opportunities. In contrast, those who attribute disparities as being internal might be more likely to frame this as a personal responsibility problem. That is, they would argue that poor individual choices are responsible for creating the disparities across groups. While social structures can limit opportunities, people should still be held personally responsible for making good choices given whatever opportunities they do have access to.
These differences could also affect our perception of inequality. Those who attribute disparities as being external might view inequality as more of a “bug” of the system. It may reflect historically unjust policies that systematically denied resources from the groups who are most in need. They may hold more negative views of capitalism, which naturally creates and reinforces inequalities across groups(Piketty and Goldhammer, 2014). On the other hand, those who attribute disparities as being internal might view inequality as more of a “feature” of the system. It may motivate individuals to make better choices for themselves through the unequal incentives provided by the invisible hand of capitalism(Smith, 1759, 1776). They may hold more positive views of capitalism as an engine for prosperity by providing the motivation for people to work hard.
The two sides might prescribe to different and opposing philosophical schools of thought. Those who attribute disparities as being external might be more likely to prescribe to Rawl’s theory of justice(Rawls, 1999), which argues that all humans before they have even been born start from an original position of equality. From this position, our principles of justice should be chosen behind a veil of ignorance, meaning that we should ignore our own personal circumstances. Since we wouldn’t know in advance whether we would be born into the least advantaged class, we would want society to institute an economic system that benefits the least advantaged. This is known as Rawl’s difference principle. However, those who attribute disparities as being internal might be more likely to prescribe to Rand’s theory of objectivism(Peikoff, 1993). She argues that reality is objective – it is what it is and not what we want it to be. She also argues that reality is independent of one’s consciousness, so we should not imagine that we all start from Rawl’s original position of equality. Therefore, we should accept reality for what it is and then make the most out of it given our own personal circumstances.
When it comes to racial disparities, we would also see a divide in philosophical perspectives. Those who attribute racial disparities as being external might be more likely to prescribe to Bell’s critical race theory, which argues that disparities are primarily driven due to unjust social structures(Delgado and Stefancic, 1998). Historical policies like slavery, Jim Crow segregation, and redlining created social inequalities that continue to persist today. They might also be proponents of Crenshaw’s intersectionality framework, which examines how overlapping identifies like race and gender could compound discrimination or privilege(Crenshaw, 1991, 1989). They might argue in favor of racially affirmative policies to correct for those past injustices. But those who attribute racial disparities as being internal might be more likely to prescribe to Sowell’s theories, which argue that individual agency contributes to the disparities(Sowell, 2019). They might argue that the disparities can be due to differences in culture, beliefs, and practices. They might also argue that racially affirmative policies could cause more harm than good, and therefore argue in favor of racially blind policies that reward people based solely on their merit and qualifications(Sowell, 2004).
This distinction between external and internal attributions can help explain the differences in preference between group fairness and individual fairness(Mitchell et al., 2021; Castelnovo et al., 2022; Verma and Rubin, 2018; Mehrabi et al., 2021; Binns, 2019). Those who attribute disparities as being external may prefer group fairness, meaning that different groups (regardless of differences in average qualifications) should receive similar outcomes on average. After all, there should be no differences between groups because they cannot be blamed for any disparities. Those who attribute disparities as being internal may prefer individual fairness, meaning that similarly qualified individuals (regardless of group membership) should receive similar treatment. That is, individuals who want better outcomes should simply work harder so that they can prove themselves through their own merit and qualifications.
Finally, the divide between internal versus external attribution forms the basis of the political and legal debate over how to close the disparities between protected groups(Sowell, 2019; Donald, 2023; Mills, 1997; Kendi, 2017; Loury, 2002). Those on the left-side of the political spectrum (e.g. liberals and progressives) are more likely to attribute disparities as being external. This would make them believe that the system and the data that it produces is fundamentally biased and unjust. They might argue for large government interventions like affirmative action or DEI to correct for social injustices. They might focus more on preventing disparate impact (also known as indirect discrimination), a form of illegal discrimination where a neutral policy may disproportionately impact one group more than another. Meanwhile, those on the right-side of the political spectrum (e.g. conservatives and libertarians) are more likely to attribute disparities as being internal. This would make them believe that the system and the data that it produces is fundamentally unbiased and just. They might argue for limited government intervention like fairness through unawareness. They might focus more on preventing disparate treatment (also known as direct discrimination), a form of illegal discrimination where individuals are being treated differently based on their membership in a protected group.
This gives us a basic philosophical framework for understanding society’s different ideological viewpoints about whether a system is biased and unfair. We will now encode these ideas for modeling our different beliefs about bias from the data generating process into several causal graphs. By understanding society’s diverse and oftentimes contradictory beliefs about the origins of bias, we can use these graphs to model different forms of discrimination. This will allow us to understand the different types of interventions people might want to employ to ensure that a decision is fair.
4. Modeling fairness as a complex system
The algorithmic fairness problem can be modeled as a complex system by merging techniques from machine learning, causal inference, and system dynamics. Each of these techniques capture different aspects of the fairness and can be seen as operating at different layers within a larger causal hierarchy. The first layer is machine learning, where we want models to make fair predictions and decisions that are compliant with society’s laws and regulations. The next layer is causal inference, where we can model different forms of bias and discrimination. This helps us to identify what interventions are necessary to make a fair decision. The final layer is system dynamics, where we can model the dynamics of the fairness problem to identify counterintuitive behavior within the system. This allows us to build mental models for how different fairness interventions can affect disparities across protected groups.
Figure 1 provides a visual representation of these three different but complementary analytical techniques. Notice that in each representation, we try to use the same variables to demonstrate that they are different ways of looking at the same problem. Examples of what each of the variables in these figures might represent are listed for several high-stakes applications in Table 1.

### v3

@@ Abstract.
Systems thinking provides us with a way to model the algorithmic fairness problem by allowing us to encode prior knowledge and assumptions about where we believe bias might exist in the data generating process. We can then encode these beliefs as a series of causal graphs, enabling us to link AI/ML systems to politics and the law. This allows us to combine techniques from machine learning, causal inference, and system dynamics in order to capture different emergent aspects of the fairness problem. We can use systems thinking to help policymakers on both sides of the political aisle to understand the complex trade-offs that exist from different types of fairness policies, providing a sociotechnical foundation for designing AI policy that is aligned to their political agendas.
@@ 1. Introduction
For several years, AI researchers have struggled with how to build AI/ML systems that are fair and aligned to society’s values. Purely technical approaches that rely on correlation alone have fallen far short of what is needed to solve this difficult problem. Systems thinking offers a novel approach to the algorithmic fairness problem by allowing us to use causality to unify complex ideas from the social sciences and computer science into a comprehensive and integrated theory. By leveraging prior knowledge and assumptions about where we believe bias might exist in the data generating process, we can help policymakers on both sides of the aisle to design AI policy that is aligned to their political agendas.
To address the algorithmic fairness problem, we need to take a sociotechnical approach. The first part of this paper will briefly cover concepts from the social sciences to address the controversial question: Why do we believe disparities exist across protected groups like race or gender? This will give us insights into the political debate around fairness and why policymakers fundamentally disagree on how to close those disparities.
We will then turn our attention to the technical aspects of the fairness problem, which forms the bulk of this paper. We will use systems thinking to show how to combine analytical techniques from machine learning, causal inference, and system dynamics to model different facets of the fairness problem. Machine learning is necessary for building AI systems that can make fair predictions. Causal inference allows us to visualize different forms of bias, showing us what interventions are needed to make a machine learning model fair. System dynamics allows us to model the bias in the data generating process, providing counterintuitive explanations as to why disparities continue to persist across protected groups. Finally, we will demonstrate how these three techniques can form the basis of a new causal hierarchy for modeling algorithmic fairness.
@@ 2. Related work
The field of algorithmic fairness is a relatively new field that has drawn significant attention since the release of a ProPublica article that claimed racial bias in COMPAS, a software tool that is widely used in the US court system to predict recidivism(Angwin et al., 2016). Since then, there have been multiple attempts to align the legal definitions of discrimination (e.g. disparate treatment and disparate impact in the US or direct discrimination and indirect discrimination in the EU) with their equivalent mathematical definitions. But these approaches, which have oftentimes relied on statistical correlation, have not been successful. Narayanan for example identified nearly two dozen fairness metrics, none of which could fully align to the legal definitions of discrimination(Narayanan, 2018).
@@
Pearl has argued for a few decades that the discrimination problem is fundamentally a causality problem(Pearl, 2001). He has argued that fairness can be modeled using causal Bayesian networks (CBNs) as a series of direct, indirect, and spurious effects. This work was later expanded upon by Plecko and Bareinboim, who proposed the idea of a standard fairness model(Plecko and Bareinboim, 2022). They argued that there exists a causal template through which entire classes of fairness models can be derived using a small number of modeling assumptions. Kusner et al. argued for the use of counterfactuals in modeling fairness(Kusner et al., 2017). Chiappa and Isaac showed that certain paths within a CBN could be labeled as fair or unfair(Chiappa and Isaac, 2019). Another attempt of building causal fairness models was made by Nilforoshan et al., but their approach had encountered limitations(Nilforoshan et al., 2022). Barocas et al. devoted a chapter in their textbook on fairness and machine learning to discuss how causality might be used to model fairness(Barocas et al., 2023).
@@
However, their causal approaches were not really sociotechnical and didn’t take full account of the social and political aspects of the fairness problem. Society has different ideological beliefs about what is fair and unfair based on different sets of assumptions about where bias is in the data generating process. Our paper addresses this issue by encoding society’s different mental models of bias into causal models using first principles. By taking into account a diversity of perspectives, we can understand the normative aspects of the fairness problem. This is necessary for linking machine learning to politics and ultimately the law.
@@
Another key difference is that their causal approaches relied solely on causal inference, which has limitations in its ability to model the data generating process. Our approach leverages systems thinking and specifically system dynamics to gain a fuller understanding of the data generating process through the modeling of feedback loops. Martin has also explored modeling AI bias using community-based system dynamics(Jr, 2023, 2021). Furthermore, we link supervised machine learning to causal inference by recognizing that a CBN can be used as a higher level abstraction of a black box ML model. By bringing machine learning, causal inference, and systems thinking together, we can model the different aspects of the algorithmic fairness problem.
@@ 3. Towards a unified theory of fairness
To build fair AI/ML systems, we begin from first principles with the strong assumption that all humans have free will(Carus, 1910). This forms the basis of moral responsibility(Talbert, 2024). If we humans have free will, then we have the agency to make our own decisions. These decisions will lead to certain outcomes, and we may be held morally responsible for those outcomes.
@@
But sociologists would note that we are never fully free to make our own decisions. This is because social structures can influence our agency(Barker, 2005). Social structures define our relationships with the various institutions of society and may include our economic systems, legal frameworks, cultural norms, and political institutions. Agency defines our ability to act independently and to freely make our own decisions. Social structures place a limit on the the choices and opportunities that are available to us humans. Those who are economically disadvantaged will encounter greater limits to their agency due to social structures compared to those who are economically advantaged.
@@
To better understand human behavior, we turn to a concept in psychology called the locus of control. This was coined by Rotter and refers to the degree of belief that people have control over the outcomes of events in their own lives(Rotter, 1966). Humans with an internal locus of control believe that they are in control of their own lives, while those with an external locus of control believe that their lives are controlled by outside factors. These concepts might also be related to whether someone adopts a growth mindset(Dweck, 2007) or a victimhood mindset(Kaufman, 2020).
@@
Psychologists provide another key concept for fairness called attribution theory, which is used by humans to interpret and assign causes for events. First proposed by Heider(Heider, 1958) and later expanded by Kelley(Kelley, 1967) and Weiner(Weiner, 1986), attribution theory examines whether attributions should be assigned to internal (dispositional) attributions or external (situational) attributions. The internal attributions may be due to behavior or personal traits, while the external attributions may be due to situational or environmental factors. This is essential for determining whether we believe an outcome is biased and therefore unfair.
@@
For example, let’s say that a student did poorly on an exam. An internal attribution might argue that the student was lazy and should have spent more time studying. In this case, the student is to blame for the poor results and thus the test was unbiased and fair. An external attribution might argue that the test was too hard or that the material was not taught well. In this case, the professor is to blame for the poor results and thus the test was biased and unfair.
To determine whether an outcome is biased and unfair, all we need to know is whether we would attribute a bad outcome as being internal or external. This makes algorithmic fairness a type of credit assignment problem. But since we usually focus on bad outcomes, this oftentimes becomes a blame assignment problem. That being said, identifying attribution is much harder in practice, as we will show later when we model fairness using system dynamics. Both internal and external factors may influence and reinforce each other, leading to vicious feedback loops. This becomes a particularly difficult problem when trying to explain disparities across protected groups like race, which we describe in greater detail in Appendix A.
This causality dilemma between attributing disparities to internal or external factors ultimately leads to the controversial and heated political debate over how best to reduce disparities across protected groups(Sowell, 2019; Donald, 2023; Mills, 1997; Kendi, 2017; Loury, 2002). Those on the political left are more likely to attribute disparities as being due to external factors, which provides the moral justification for strong government interventions like affirmative action or DEI. The left also tends to focus more on legal forms of discrimination like disparate impact or indirect discrimination. Those on the political right are more likely to attribute disparities as being due to internal factors, which provides the moral justification for limited government intervention like fairness through unawareness. The right also tends to focus more on legal forms of discrimination like disparate treatment or direct discrimination.
We will use systems thinking to encode these different ideological beliefs into a series of causal graphs. This will help us to understand what types of interventions policymakers on both sides of the political aisle might want to make to ensure that a machine learning model is fair. Finally, this could give us insights into how to build machine learning models that are compliant with antidiscrimination laws.
4. Modeling fairness as a machine learning problem
Before we build a causal representation of the fairness problem, we will start by modeling fairness as a supervised machine learning problem using a simple block diagram as shown in Figure 1. In the center is a black box machine learning model that receives data $X$ as input and returns decision $D$ as output. The decision $D$ causes an outcome $Y$ that could potentially be recorded as future data $X$, thus creating a feedback loop. The key question that we need to address is whether the protected attribute $A$ should be used in a machine learning model and if so, how?
[…]
Figure 1. Block diagram representation of algorithmic fairness
[…]
A block diagram of a black box ML model receiving data $X$ as input and returning decision $D$ as output. Decision $D$ causes an outcome $Y$ which can become future data $X$. A protected attribute $A$ is being fed into the model.
5. Modeling fairness as a causal inference problem
We can transform the block diagram from Figure 1 into a causal Bayesian network (CBN) as shown in Figure 2. CBNs were pioneered by Pearl(Pearl, 2000; Pearl and Mackenzie, 2018; Pearl et al., 2016) and allow us to symbolically encode knowledge and assumptions about causal effects between variables, which are shown in this graph as nodes. The first node of the CBN represents a protected attribute $A$, which has a direct effect on a mediator $W$ ($A\rightarrow W$). The mediator $W$ has a direct effect on an outcome $Y$ ($W\rightarrow Y$). We assume that the mediator $W$ completely explains away the causal effect between the protected attribute $A$ and the outcome $Y$, resulting in conditional independence ($Y\perp\!\!\!\perp A|W$). However, the mediator $W$ might be considered biased by some people if they attribute the disparities across protected groups as being external. The mediator $W$ is filled white because it is latent, but it can still be partially measured as data $X$ ($W\rightarrow X$). Data $X$ is then fed into a machine learning model that is embedded inside the node for decision $D$ ($X\rightarrow D$). Finally, the decision $D$ also has a direct effect on an outcome $Y$ ($D\rightarrow Y$). CBNs are represented by directed acyclic graphs (DAGs), which means that we cannot use them to model feedback loops. Examples of what each of these variables could represent are shown in Table 1. For each of these applications, the protected attribute $A$ could represent race or gender for example.
[…]
A causal Bayesian network where the protected attribute $A$ causes mediator $W$, which in turn causes an outcome $Y$ and is recorded as data $X$. The data $X$ is used to make a decision $D$ which can also affect outcome $Y$.